OpenStudy (anonymous):

16. Solve the equation step by step. List each step, along with the property you used during that step. -5*1*11*4

OpenStudy (anonymous):

I was thinking that it would be -5*1*11*4= -5*1=-5 -5*11=-55 -55*4=-220 is that right

OpenStudy (anonymous):

wrong subject bud